OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (anonymous):

how am i supposed to answer this if i dont know the common ratio ?

OpenStudy (whpalmer4):

If you have two terms in the sequence, \(n_1\) and \(n_2\), you know that the common ratio is \(a = \frac{n_2}{n_1}\) right?

OpenStudy (anonymous):

noo, so its 8/10 ?

OpenStudy (whpalmer4):

What if you had two terms, but they were \(n_1\) and \(n_3\)? Then you know that \(n_2=n_1*a\) and \(n_3= n_2*a\) so \(n_3=(n_1*a)*a = n_1*a^2\)

OpenStudy (whpalmer4):

If you knew the 1st and 4th terms, then \(n_4 = n_1*a*a*a = n_1*a^3\) can you see the pattern developing here?

OpenStudy (anonymous):

this is confusing :(. is that how to find the common ratio ?

OpenStudy (whpalmer4):

Yes. If you know terms 1 and n of a geometric sequence, then term n = term 1 * common ratio to the (n-1) power. Let's try an example. Say the common ratio is 2. We know that the first term is 3, and we want to find the 3rd term. Well, that's going to be 3*2*2 = 12, right? Suppose we instead knew that we had a 3 term sequence, the first term was 3, and the 3rd term was 12. We could find the common ratio by dividing 12/3 = 4. 4 represents the product of all of the common factor multiplications to go from the 1st term to the 3rd term. That's 2 terms, so 4 must be the common ratio squared, right? and indeed, 4 = 2*2

OpenStudy (whpalmer4):

For your problem, we have 8 terms, the first is 10, and the 8th is 781250. If we divide the 8th by the 1st, we get 78125. Our common ratio to the 7th power (8-1) = 78125. We could just try some numbers: 2^7 = 128, not big enough. 10^7 = 10000000, too big. How about 5^7? What do you know, 5^7 = 78125! So our common ratio is 5.

OpenStudy (whpalmer4):

Given that the ratio ended in 5, 5 was actually a pretty obvious candidate to try first, as all numbers ending in 5 or 0 are divisible by 5...

OpenStudy (anonymous):

This is soooo hard :/

OpenStudy (whpalmer4):

It's a bit tricky at first, I agree.

OpenStudy (anonymous):

for geometric sequence a term is given by T=a(r^n-1) where n is the number of term to be found in this case for last term T=781250 n=8 a=10 plug in values and solve for r

OpenStudy (anonymous):

and its T=a(r^(n-1))

OpenStudy (whpalmer4):

@lesha here's what you see for the first 8 terms for common ratios from 2 through 7: 2: {10, 20, 40, 80, 160, 320, 640, 1280} 3: {10, 30, 90, 270, 810, 2430, 7290, 21870} 4: {10, 40, 160, 640, 2560, 10240, 40960, 163840} 5: {10, 50, 250, 1250, 6250, 31250, 156250, 781250} 6: {10, 60, 360, 2160, 12960, 77760, 466560, 2799360} 7: {10, 70, 490, 3430, 24010, 168070, 1176490, 8235430}

OpenStudy (whpalmer4):

You can see that a common ratio of 5 is the one that gets you from 10 to 781250 in 8 terms.